OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (texaschic101):

do you know the slope of the equation x - 2y = 6 ?

OpenStudy (anonymous):

i have x-2y-10

OpenStudy (texaschic101):

to find the slope, we will put the equation in y = mx + b form, and the number in the m position is the slope. x - 2y = 6 -- subtract x from both sides -2y = -x + 6 -- divide by -2 y = 1/2x - 3 the number in the m position is 1/2, therefore, your slope is 1/2. A parallel line will have the same slope. Now we will use y = mx + b slope(m) = 1/2 (-2,4) x = -2 and y = 4 now we sub and solve for b 4 = 1/2(-2) + b 4 = - 1 + b 4 + 1 = b 5 = b so your parallel equation is : y = 1/2x + 5 but you need it in standard form... y = 1/2x + 5 --- multiply by 2 2y = x + 10 -- subtract x -x + 2y = 10 -- multiply by -1 to make x positive x - 2y = -10

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)